# Gram-like matrix preserving extensions of noncommutative polynomials to sum of Hermitian Squares

### v1

@@ Abstract.
Given a nonnegative noncommutative polynomial $f$, equivalently a sum of Hermitian squares (SOHS), there exists a positive semidefinite Gram matrix that encrypts all essential information of $f$. There are no available methods for extending a noncommutative polynomial to a SOHS keeping the Gram matrices unperturbed. As a remedy, we introduce an equally significant notion of Gram-like matrices and provide linear algebraic techniques to get the desired extensions. We further use positive semidefinite completion problem to get SOHS and provide criteria in terms of chordal graphs and 2-regular projective algebraic sets.
@@ 1. Introduction
Hilbert’s 17th problem states whether every multivariate polynomial that takes only nonnegative values over the reals can be represented as sum of squares (SOS) of polynomials. Hilbert, in 1888, proved that it is true only for quadratic forms, binary forms and ternary quatrics in the commutative case. In the noncommutative setup, Helton proved that any polynomial is nonnegative if and only if it is a sum of Hermitian squares (SOHS). The SOS and SOHS techniques are heavily used for polynomial optimization and convex optimization, in particular, in semidefinite programming (SDP). Many of the SOHS techniques are taken from real algebraic geometry and inspired from functional analysis.
@@
The SOHS techniques apply across many different disciplines including control theory, probability and statistics. Out of many directions, hierarchy techniques are still very active areas of research in quantum information theory (cf. [12] and [22]) and have been proven useful in discriminating the quantum states. A large portion of the theory for noncommutative polynomials is being developed by Helton, McCullough, Klep and their collaborators (cf. [10], [17], [18], [20] and [21]). The book “Optimization of Polynomials in Non-commuting Variables” (cf. [9]) by Burgdorf, Klep and Povh is a beautiful introduction to the subject. Pablo Parillo has used SOS techniques in different SDP to produce different control theory oriented applications (cf. [23]). Blekherman and his collaborators’ work on SOS techniques provide fruitful insight about the fundamental geometric properties of some projective algebraic sets (cf. [8], [6] and [7]). Motivated by the theory of noncommutative polynomials, noncommutative functions have been developed over noncommutative sets and many researchers are working on the subject (cf. [1] and [25]).
This paper is aimed at a large audience with minimal background in the noncommuatative polynomials. Most of the concepts are self-contained. One of the key aspects that the reader will take away from this paper is the technique to produce ample SOHS. Moreover, these techniques will be fruitful in SDP and its applications.
This article blends different areas, such as linear algebra, graph theory, algebraic geometry and commutative algebra, with a focus on SOHS techniques. Given a noncommutative polynomial parts of which are SOHS, it is always possible to find a SOHS extension $\tilde{f}$ by adding suitable terms with $f$. But, in general, a Gram matrix of such $\tilde{f}$ doesn’t preserve the Gram matrices of the SOHS parts of $f$. In this paper, we define a notion called Gram-like matrix, which is equally significant to, but more flexible than, a Gram matrix. Exploiting the obtained flexibility and Sylvester’s criteria, we provide techniques to preserve any Gram-like matrix of the SOHS part of some well behaved $f$ inside a Gram-like matrix of $\tilde{f}$ as a leading principal submatrix. Furthermore, we prove that a SOHS completion $\tilde{f}$ of a quasi SOHS polynomial $f$ is possible if the specification graph and the subspace arrangement of $f$ is chordal and $2$-regular respectively. In that case, a Gram-like matrix of $\tilde{f}$ contains Gram-like matrices of the SOHS parts of $f$ as its principal submatrices. In fact, we show that this SOHS completion is an alternative elucidation of the celebrated positive semidefinite completion problem.
@@
We now go through the chronology of the paper in a bit more detail. In Section 2, we recall some prerequisites regarding noncommutative polynomials. We discuss the notion of symmetric polynomials, SOHS polynomials and some basic properties of positive definite and positive semidefinite matrices. Following that, we rigorously discuss the notion of a Gram matrix and monomial vector associated to a symmetric polynomial and SOHS. We observe that depending on the polynomial, the size of the associated Gram matrix can be reduced without affecting the defining properties. Moreover, one can be carefree about the order of the constituent monomials of the monomial vector. We point out these drawbacks of a Gram matrix through several examples (cf. $(2)$ of Remark 2.13 and Example 2.14). We sort these things out formally (cf. Theorem 3.5 and Theorem 3.6) in Section 3 and refer to the associated matrix as Gram-like matrix (cf. Definition 3.7 and Definition 3.8), which is analogous to, but more flexible than, Gram matrix (cf. Remark 3.9 for a comparison between a Gram matrix and a Gram-like matrix). Moreover, it keeps the soul of a Gram matrix intact. We then ask the following question :
@@
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial.
  1. (a)
    What is the procedure to get a SOHS extension $\tilde{f}$ of $f$?
  2. (b)
    If $f$ has some SOHS parts, can it be extended to a SOHS $\tilde{f}$ keeping the gram-like matrices of the the SOHS parts of $f$ intact?
Theorem 3.14 answers part $(a)$ of Question 1, whereas Theorem 3.16, Theorem 3.29 and Theorem 3.40 answers part $(b)$ of Question 1.
We first answer this question in the most trivial case, namely, adding a positive constant to a SOHS gives a SOHS (cf. Theorem 3.10). But we do so by exploiting the flexibility of a Gram-like matrix. Followed by that, we prove the first main theorem of this article (cf. Theorem 3.16).
Theorem 1.1.
Let $f\in\mathbb{R}\langle\underline{X}\rangle$ be a noncommutative polynomial, without any constant term, which can be expressed as a SOHS plus $r$ many terms $a_{1}\zeta_{1},a_{2}\zeta_{2},\ldots,a_{r}\zeta_{r}$. Assume that ${\rm rc}(\eta,i)\neq\zeta_{j}$, $1\leq j\leq r$, for any $i\geq 0$ and for any monomial $\eta$ of the SOHS part of $f$. Then
@@
    It is always possible to obtain a SOHS $\widetilde{f}$ by adding at most $12r+$ 1 terms with $f$ such that there exists a positive semidefinite Gram-like matrix of $\widetilde{f}$ which contains a positive semidefinite Gram-like matrix of the SOHS part of $f$ as a leading principal submatrix.
@@
    In fact, infinitely many such $\widetilde{f}$ can be obtained.
We further observe that one can even control the degree of $\widetilde{f}$ in the aforementioned scenario (cf. Theorem 3.18).
In subsection 3.3, we deal with Question 1 via graph theory. In fact, we look at positive semidefinite completion problem in the context of completion of a quasi SOHS polynomial (cf. Definition 3.28 for details). We prove the following theorem which describes a pattern made out of the known coefficients of a quasi SOHS polynomial that ensures a SOHS completion of the same (cf. Theorem 3.29). The pattern is given in terms of a graph associated to a partial Gram-like matrix of the given quasi SOHS polynomial.
Theorem 1.2.
@@
Theorem 1.3.
  1. (1)
  2. (2)
  3. (3)
We provide some examples to display simultaneous applications of Theorem 3.29 and Theorem 3.40. In these examples we use Macculay2 software to calculate regularity via Betti tables and free resolutions.
@@ 2.2. Some properties of positive definite and semidefinite matrices
There are many different equivalent definitions for a symmetric matrix to be positive semidefinite. Here we present one with the matrix minors. A minor of a square matrix is the determinant of a submatrix obtained by deleting same number of rows and columns. A principal submatrix of a square matrix is the one where the indices of the deleted rows and the indices of the deleted columns are same. A principal minor is the determinant of a principal submatrix. A leading principal submatrix is a principal submatrix where index of the one deleted row (or column respectively) implies all the rows (or column respectively) with higher indices are deleted. In other words, the leading principal submatrices are square matrices sitting at the left corners of the actual matrix. Determinant of a leading principal matrix is called a leading principal minor.
@@
Theorem 2.4 (Sylvester’s criterion for positive definite matrices).
@@
Analogous version doesn’t hold for a positive semidefinite matrix, i.e, nonnegativity of all leading principal minors doesn’t guarantee the positive semidefiniteness of a real symmetric matrix. For example, both the leading principal minors of the real symmetric matrix
$\begin{pmatrix}0&0\\ 0&-1\end{pmatrix}$
are 00. But the matrix clearly is not positive semidefinite as its eigenvalues are 00 and $-1$ (cf. [3, §1.2, p. 7]). Therefore, we need some extra conditions in case of positive semidefinite matrices.
Theorem 2.5 (Generalized Sylvester’s criterion for positive semidefinite matrices).
@@
We end this subsection by recalling one more well-known property of a positive semidefinite matrix. We present a rudimentary argument, which is available in the literature, for the reader not familiar with this result.
@@
Follows from Theorem 2.5. Indeed, suppose the contrary. Then there will be some $1\leq j\leq k$ such that $(i,j)$-th entry, and hence $(j,i)$-th entry of $A$ is nonzero and both are equal. Then, the determinant of the $2\times 2$ principal submatrix of $A$ formed by $(i,i)$-th entry, $(i,j)$-th entry, $(j,i)$-th entry and $(j,j)$-th entry is negative, which contradicts Theorem 2.5. ∎
@@ 2.3. Gram Matrix of a symmetric polynomial and a SOHS
$\begin{split}&a_{i},b_{ij}\in\mathbb{R}\;;\\ &\zeta_{i}\in\mathbb{R}\langle\underline{X}\rangle,\;{\rm deg}(\zeta_{i})\leq d\text{\;for\;all\;}1\leq i\leq k\;;\\ &s(d,n)=\sum_{k=0}^{d}n^{k}\;.\end{split}$ (2)
@@
    However, any symmetric homogeneous polynomial of degree $2d$ has a unique representation as all terms of such a polynomial are of degree $2d$ and hence can be written uniquely in terms of product of two degree $d$ monomials.
  3. (3)
    The moment we fix a representation of a degree $2d$ symmetric noncommutative polynomial, associated Gram matrix and monomial vector get fixed. Therefore, any symmetric homogeneous polynomial of degree $2d$ has a unique Gram matrix, as mentioned in [9, Proposition 1.19, p. 7].
@@
  1. (1)
    At this point it is worthwhile to mention that to be able to write a SOHS polynomial $f$ as $f=W_{s(d,n)}^{\ast}\;G\;W_{s(d,n)}$ for some symmetric matrix $G$ doesn’t guarantee the nonnegativity of the matrix $G$. For example, we consider the polynomial (cf. [4, Example 3.4, p. 3])
    $f=1+2X_{1}+X_{1}^{2}+X_{1}X_{2}^{2}+2X_{2}^{2}+X_{2}^{2}X_{1}+X_{2}X_{1}^{2}X_{2}+X_{2}^{4}.$
    Then $f$ is an SOHS as it can be decomposed as
    $f=(1+X_{1}+X_{2}^{2})^{\ast}(1+X_{1}+X_{2}^{2})+(X_{1}X_{2})^{\ast}(X_{1}X_{2}).$
    Taking
    $G_{f}=\begin{pmatrix}1&1&0&0&0&0&0\\ 1&1&0&0&0&0&1\\ 0&0&0&0&0&0&0\\ 0&0&0&1&0&0&0\\ 0&0&0&0&2&0&0\\ 0&0&0&0&0&0&0\\ 0&1&0&0&0&0&1\end{pmatrix}\;,$
    we have $f=W_{7}^{\ast}\;G_{f}\;W_{7}$, where $W_{7}$ is given as
    $W_{7}=\begin{pmatrix}1&X_{1}&X_{1}^{2}&X_{1}X_{2}&X_{2}&X_{2}X_{1}&X_{2}^{2}\end{pmatrix}^{t}\;.$
    It can be noted that the symmetric matrix $G$ is not positive semidefinite as determinant of the principal sub matrix
    $\begin{pmatrix}1&1&0&0\\ 1&1&0&1\\ 0&0&1&0\\ 0&1&0&1\end{pmatrix}\;,$
    obtained from $G_{f}$ by deleting third, fifth and sixth row and column, is $-1$. One can look at [20, Example 2.2, p. 742] or [9, Example 1.18, p. 6-7] for another such example.
  2. (2)
    To express a symmetric polynomial of degree $2d$ in the form as in (3), it is not always necessary to take all the monomials of degree at most $d$ in $W_{s(d,n)}$. For example, consider the polynomial $f$ as in (1). Then for
@@
    we have $f=W_{5}^{\ast}\;G\;W_{5}$.
As the choice of the monomial vector $W_{s(d,n)}$ is very strict in Theorem 2.9 and Theorem 2.12, there are several techniques to throw out the unnecessary monomials and making the matrix of smaller size. For example Newton chip method, augmented Newton chip method etc (cf. [9, Section 2.3 & 2.4]) has been used to reduce the size of the column vector $W_{s(d,n)}$, as in (5), and hence the size of matrix $G_{f}$. Moreover, while putting the monomials in $W_{s(d,n)}$, usually lexicographic order is followed. But as long as enough monomials is being collected in the monomial vector to express a given SOHS $f$, the order doesn’t matter. We demonstrate this through the following example.
@@
Take $f=X_{1}^{2}+4X_{1}X_{2}+4X_{2}X_{1}+16X_{2}^{2}$. Then, for $(W^{\prime}_{2})^{\ast}=\begin{pmatrix}X_{1}&X_{2}\end{pmatrix}$, we have $G^{\prime}=\begin{pmatrix}1&4\\ 4&16\end{pmatrix}$ such that $f=(W^{\prime}_{2})^{\ast}\;G^{\prime}\;W_{2}^{\prime}$. Now let’s change the order of the monomials. Take, $(W^{\prime\prime}_{2})^{\ast}=\begin{pmatrix}X_{2}&X_{1}\end{pmatrix}$, we have $G^{\prime\prime}=\begin{pmatrix}16&4\\ 4&1\end{pmatrix}$ such that $f=(W^{\prime\prime}_{2})^{\ast}\;G^{\prime\prime}\;W_{2}^{\prime\prime}$.
It can be noted that the matrix $G^{\prime\prime}$ is obtained from the positive semidefinite matrix $G^{\prime}$ by interchanging its rows and columns and still is remaining a positive semidefinite matrix. This is what is allowing us not to bother about the order of the monomials in $W_{2}$. Lemma 2.6 says that this is not at all surprising.
It is evident from part $(2)$ of Remark 2.13 and Example 2.14 that to be able to reduce the size of the monomial vector (and hence that of a Gram matrix) whenever possible and write the monomial entries in that monomial vector in any arbitrary order, while expressing a symmetric or SOHS polynomial in the required form, puts us in a resilient situation. Therefore, we formalize all of this in the upcoming section after putting things in a particular perspective.

### v2

# Gram-like matrix preserving extensions and completions of noncommutative polynomials

@@ Abstract.
Given a positive noncommutative polynomial $f$, equivalently a sum of Hermitian squares (SOHS), there exists a positive semidefinite Gram matrix that encrypts all the structural essence of $f$. There are no available methods for extending a noncommutative polynomial to a SOHS keeping the Gram matrices unperturbed. As a remedy, we introduce an equally significant notion of Gram-like matrices and provide linear algebraic techniques to get the desired extensions. We further use positive semidefinite completion problem to get SOHS and provide criteria in terms of chordal graphs and 2-regular projective algebraic sets.
@@ 1. Introduction
Hilbert’s 17th problem states whether every multivariate polynomial that takes only nonnegative values over the reals can be represented as sum of squares (SOS) of polynomials. Hilbert, in 1888, proved that it is true only for quadratic forms, binary forms and ternary quatrics in the commutative case. In the noncommutative setup, Helton proved that any polynomial is positive if and only if it is a sum of Hermitian squares (SOHS). The SOS and SOHS techniques are heavily used for polynomial optimization and convex optimization, in particular, in semidefinite programming (SDP). Many of the SOHS techniques are taken from real algebraic geometry and inspired from functional analysis.
@@
The SOHS techniques apply across many different disciplines including control theory, probability and statistics. Out of many directions, hierarchy techniques are still very active areas of research in quantum information theory (cf. [12] and [23]) and have been proven useful in discriminating the quantum states. A large portion of the theory for noncommutative polynomials is being developed by Helton, McCullough, Klep and their collaborators (cf. [10], [17], [18], [20] and [21]). The book “Optimization of Polynomials in Non-commuting Variables” (cf. [9]) by Burgdorf, Klep and Povh is a beautiful introduction to the subject. Pablo Parillo has used SOS techniques in different SDP to produce different control theory oriented applications (cf. [24]). The work of Blekherman and his collaborators on SOS techniques provide fruitful insight about the fundamental geometric properties of some projective algebraic sets (cf. [8], [6] and [7]). Motivated by the theory of noncommutative polynomials, noncommutative functions have been developed over noncommutative sets and many researchers are working on the subject (cf. [1] and [26]).
Our main contributions are threefold. First, we provide a linear algebraic construction to extend noncommutative polynomials to SOHS, by adding necessary terms, that ensures the Gram-like matrix of the SOHS part is a principal submatrix of the extension’s matrix (cf. Theorem 3.24). Second, we leverage chordal graph theory to characterize when a partially specified noncommutative polynomial can be completed to a SOHS, linking matrix completion to SOHS completion (cf. Theorem 3.35). Third, we connect these completions to projective algebraic sets, showing that their 2-regularity guarantees SOHS completions (cf. Theorem 3.47). These results blend linear algebra, graph theory, and algebraic geometry. One of the key aspects that the reader will take away from this paper is some techniques to produce ample SOHS. Moreover, these techniques will be fruitful in polynomial optimization.
@@
We now go through the chronology of the paper in a bit more detail. In Section 2, we recall some prerequisites and discuss the notion of a Gram matrix & monomial vector associated to a symmetric polynomial & SOHS. We introduce the notion of a Gram-like matrix which reduces the size of Gram matrix and disregards the order of the constituent monomials in the monomial vector (cf. Remark 3.9 for a detailed comparison between a Gram matrix and a Gram-like matrix). Moreover, it keeps the soul of a Gram matrix intact.
The following question captures the central theme of this paper:
@@
Given a noncommutative polynomial $f\in\mathbb{R}\langle\underline{X}\rangle$ that includes SOHS components, can we construct a SOHS polynomial $\widetilde{f}\in\mathbb{R}\langle\underline{X}\rangle$ that extends (or completes) $f$ such that the positive semidefinite Gram-like matrices of the SOHS components are principal submatrices of the positive semidefinite Gram-like matrix of $\tilde{f}$?
Theorem 3.24, Theorem 3.35 and Theorem 3.47 answer Question 1 in three different set up in three consecutive subsections.
@@
    In subsection 3.2, assuming that there exists some $\widetilde{f}$ satisfying the properties as in Question 1, we first note that the structure of a positive semidefinite Gram-like matrix $G_{\widetilde{f}}$ of such $\widetilde{f}$ would be of the form $\begin{pmatrix}G_{h}&A\\ A^{t}&B\end{pmatrix}$, where $G_{h}$ is some positive semidefinite Gram-like matrix of $h$ (cf. Lemma 3.14). Then we provide a way to produce a positive semidefinite matrix out of a partially specified symmetric matrix of the aforementioned structure only by controlling the diagonal elements of $B$ (cf. Lemma 3.17. Exploiting these two results, we further provide some necessary and sufficient conditions in terms of a representation of a SOHS part $h$ of $f$ and monomial decompositions of $f-h$ (cf. Theorem 3.19).
    We provide necessary and sufficient condition for $f=h+\sum_{j}d_{j}\delta_{j}$, with $h$ SOHS and $\delta_{j}$ are not hermitian squares, to have a SOHS extension $\tilde{f}$. If for any decomposition of $\delta_{j}=l_{k}r_{k}$, $l_{k}$ and $r_{k}$ simultaneously do not belong to the set of monomial vectors of a Gram-like matrix $G_{h}$ and if $G_{h}$ is positive definite, we can always extend $f$ to a Gram-like matrix preserving $\tilde{f}$. But if $G_{h}$ is psd but not pd, we need to have additional conditions, namely, the column space of $A$ should be sitting inside the column space of $G_{h}$. We see that the problem boils down to choosing diagonal elements of lower diagonal block of a $2\times 2$ block matrix, so that we can make the block matrix positive semidefinite.
    This, in turn, helps us to find noncommutative polynomials $f$ that do not admit any such extension $\widetilde{f}$ (cf. Example 3.13). We then provide necessary and sufficient conditions for the existence of such $\widetilde{f}$, when $G_{\widetilde{f}}$ takes a special form, namely as in (11) (cf. Theorem 3.22). This characterisation helps us to produce a plenty of SOHS polynomials such that their difference is also a SOHS, which is not true in general (cf. Remark 3.23).
    Apart from these existential criteria of such $\widetilde{f}$ in Theorem 3.19 and Theorem 3.22, we also provide a rigorous construction of $G_{\widetilde{f}}$ (and hence $\widetilde{f}$) for a class of noncommutative polynomials $f=h+\sum_{j=1}^{r}a_{j}\zeta_{j}$, $h$ being SOHS. We do so whenever ${\rm rc}(\eta,i)\neq\zeta_{j}$, $1\leq j\leq r$, or equivalently, whenever the monomials $\zeta_{j}$ are not equal to any right end part of any monomials $\eta$ of $h$ (cf. Theorem 3.24).
@@
    In subsection 3.3, we deal with Question 1 via graph theory. In fact, we look at positive semidefinite completion problem in the context of completion of a quasi SOHS polynomial (cf. Definition 3.34 for details). We prove the following theorem which describes a pattern made out of the known coefficients of a quasi SOHS polynomial that ensures a SOHS completion of the same (cf. Theorem 3.35). The pattern is given in terms of a graph associated to a partial Gram-like matrix of the given quasi SOHS polynomial. To be specific, we prove the following :
@@
    Let $f\in{\rm Sym}\;\mathbb{R}\langle\underline{X}\rangle$ be a quasi SOHS polynomial w.r.t a representation $\mathbf{R}_{f}$, as in (8), along with conditions as in (9). Then any such $f$ has a SOHS completion $\overline{f}^{\;G_{f},\;W_{k}}$ w.r.t the partial positive semidefinite Gram-like matrix $G_{f}$ associated with $\mathbf{R}_{f}$ if and only if the subspace arrangement $V_{G_{f}}(\subseteq\mathbb{P}^{k-1})$ of $G_{f}$ is $2$-regular.
    We provide some examples to display simultaneous applications of Theorem 3.35 and Theorem 3.47. In these examples we use Macaulay2 software to calculate regularity via Betti tables and free resolutions.
@@ 2.1. Noncommutative polynomial and SOHS
A noncommunicative polynomial $f\in\mathbb{R}\langle X\rangle$ is positive if $f(A_{1},\dots,A_{n})$ is positive semidefinite for all tuples $(A_{1},\dots,A_{n})$ of real symmetric matrices of the same order.
Definition 2.2.
@@
Remark 2.3.
A noncommutaive polynomial is positive if and only if it is SOHS. See [16] for Helton’s proof of the result and see [22] for McCullough and Putinar’s alternative proof.
@@ 2.2. Some properties of positive definite and semidefinite matrices
Lemma 2.6 (Sylvester’s criterion for positive definite matrices).
@@
Lemma 2.7 (Generalized Sylvester’s criterion for positive semidefinite matrices).
@@
We recall another well-known property of a positive semidefinite matrix. We present a rudimentary argument, which is available in the literature, for the reader not familiar with this result.
@@
Follows from Lemma 2.7. Indeed, suppose the contrary. Then there will be some $1\leq j\leq k$ such that $(i,j)$-th entry, and hence $(j,i)$-th entry of $A$ is nonzero and both are equal. Then, the determinant of the $2\times 2$ principal submatrix of $A$ formed by $(i,i)$-th entry, $(i,j)$-th entry, $(j,i)$-th entry and $(j,j)$-th entry is negative, which contradicts Lemma 2.7. ∎
We end this subsection by recalling a couple more well-known characterizations of a positive semidefinite matrix in terms of a given positive semidefinite principal minor and its Schur complement.
Lemma 2.11.
Let $M$ be a symmetric matrix with real entries of the following form
$M=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}A\end{matrix}&B\\ \hline\cr B^{t}&\begin{matrix}C\end{matrix}\end{array}\right)\;,$
where $A$ and $C$ are square matrices. Then $M$ is positive semidefinite (respectively positive definite) if and only if $A$ is positive definite and $M/A(:=C-B^{t}A^{-1}B)$ is positive semidefinite (respectively positive definite).
Proof.
See [27, Theorem 1.12, p. 34]. ∎
In fact, this lemma has a generalisation. By $A^{-}$ we denote a generalised inverse of a given square matrix $A$, i.e, $AA^{-}A=A$. Given a matrix $B$, not necessarily square, by $\mathcal{C}(B)$ we denote its column space.
Lemma 2.12.
Let $M$ be a symmetric matrix with real entries of the following form
$M=\left(\begin{array}[]{@{}c|c@{}}\begin{matrix}A\end{matrix}&B\\ \hline\cr B^{t}&\begin{matrix}C\end{matrix}\end{array}\right)\;,$
where $A$ and $C$ are square matrices. Then $M$ is positive semidefinite if and only if $A$ is positive semidefinite, $\mathcal{C}(B)\subseteq\mathcal{C}(A)$, and $M/A(:=C-B^{t}A^{-}B)$ is positive semidefinite.
Proof.
See [27, Theorem 1.20, p. 44]. ∎
@@ 2.3. Gram Matrix of a symmetric polynomial and a SOHS
$\begin{split}&s(d,n)=\sum_{k=0}^{d}n^{k}\;,\\ &a_{i},b_{ij}\in\mathbb{R}\;,\zeta_{i}\in\mathbb{R}\langle\underline{X}\rangle,\;{\rm deg}(\zeta_{i})\leq d,\;\zeta_{i}\neq\zeta_{j}\text{\;for\;all\;}1\leq i<j\leq s(d,n)\;.\end{split}$ (2)
@@
    However, the moment we fix a representation of a degree $2d$ symmetric noncommutative polynomial, associated Gram matrix and monomial vector get fixed.
@@
To express a symmetric polynomial of degree $2d$ in the form as in (3), it is not always necessary to take all monomials of degree at most $d$ in $W_{s(d,n)}$. For example, consider the polynomial $f=1+2X_{1}+X_{1}^{2}+X_{1}X_{2}^{2}+2X_{2}^{2}+X_{2}^{2}X_{1}+X_{2}X_{1}^{2}X_{2}+X_{2}^{4}$. Then for
@@
we have $f=W_{5}^{\ast}\;G\;W_{5}$. But the following Gram matrix of $f$
$G_{f}=\begin{pmatrix}1&1&0&0&0&0&0\\ 1&1&0&0&0&0&1\\ 0&0&0&0&0&0&0\\ 0&0&0&1&0&0&0\\ 0&0&0&0&2&0&0\\ 0&0&0&0&0&0&0\\ 0&1&0&0&0&0&1\end{pmatrix}$
is of much bigger size.
As the choice of the monomial vector $W_{s(d,n)}$ is very strict in Theorem 2.13 and Theorem 2.16, there are several techniques to throw out the unnecessary monomials and making the matrix of smaller size. For example Newton chip method, augmented Newton chip method etc (cf. [9, Section 2.3 & 2.4]) has been used to reduce the size of the column vector $W_{s(d,n)}$, as in (5), and hence the size of matrix $G_{f}$. Moreover, while putting the monomials in $W_{s(d,n)}$, usually lexicographic order is followed. But as long as enough monomials is being collected in the monomial vector to express a given SOHS $f$, Lemma 2.8 says that the order doesn’t matter. We demonstrate this through the following example.
@@
Consider $f=X_{1}^{2}+4X_{1}X_{2}+4X_{2}X_{1}+16X_{2}^{2}$. Then, for $(W^{\prime}_{2})^{\ast}=\begin{pmatrix}X_{1}&X_{2}\end{pmatrix}$, $(W^{\prime\prime}_{2})^{\ast}=\begin{pmatrix}X_{2}&X_{1}\end{pmatrix}$, $G^{\prime}=\begin{pmatrix}1&4\\ 4&16\end{pmatrix}$ and $G^{\prime\prime}=\begin{pmatrix}16&4\\ 4&1\end{pmatrix}$, we have $(W^{\prime}_{2})^{\ast}\;G^{\prime}\;W_{2}^{\prime}=f=(W^{\prime\prime}_{2})^{\ast}\;G^{\prime\prime}\;W_{2}^{\prime\prime}$.
It is evident from Remark 2.17 and Example 2.18 that to be able to reduce the size of the monomial vector (and hence that of a Gram matrix) whenever possible and write the monomial entries in that monomial vector in any arbitrary order, while expressing a symmetric or SOHS polynomial in the required form, puts us in a resilient situation. Therefore, we formalize all of this in the upcoming section after putting things in a particular perspective.